OpenStudy (anonymous):

Solve 7 + x - 15 = -2 2/3

OpenStudy (texaschic101):

first combine like terms 7 + x - 15 = -2 2/3 -8 + x = -2 2/3 (add 8 to both sides) x = -2 2/3 + 8 (turn your mixed fraction to an improper fraction) x = - 8/3 + 8 (common denominator is 3) x = -8/3 + 24/3 x = 16/3 or 5 1/3 check... -8 + x = 8/3 -8 + 16/3 = - 8/3 -24/3 + 16/3 = -8/3 -8/3 = -8/3 (correct) x = 16/3 or 5 1/3 any questions ?